Biochemistry Question After visualizing VCTN amounts following question 4, and calculating the average VCTN production per timepoint, you obtain th
Answer #1
[SPBRK] mM VCTN umol/min 1/[SPBRK] 1/mM 1/VCTN (min/umol)
2 6.00 0.5 0.167
4 8.19 0.25 0.122
6 9.01 0.167 0.111
8 11.68 0.125 0.0856
10 11.79 0.1 0.0848
13 12.12 0.0769 0.0825
16 15.06 0.0625 0.0664
19 14.62 0.0526 0.0684
23 14.65 0.0435 0.0683
27 15.41 0.0370 0.0649

LINEVEAVER BURK PLOT iv 1/VCTN (min/umol) y = 0.225x + 0.060 R = 0.961 00 1/VCTN 1/[SPBRK (min/um ] 1/mM ol) 0.51 0.167 0.25

Lineweaver Burk equation : 1/Vo = (Km/Vmax) (1/[S] + 1/Vmax

the above equation is comparable to y = mx+c

slope, m = Km/Vmax and Y-intercept = 1/Vmax

the equation of above straight line graph : y = 0.225x+0.060

slope, m = Km/Vmax = 0.225

Y-intercept, c = 1/Vmax = 0.060

Vmax = 1/0.060 = 16.7 uM/min ( 3 significant figures)

Km/Vmax = 0.225

Km/16.67 = 0.225

Km = 0.225 x 16.67 = 3.8 mM ( 2 significant figures)